OpenStudy (anonymous):

What is the restriction on the product of quantity x squared plus 5 x plus 4 divided by quantity x squared plus 3 x minus 4 multiplied by quantity x squared plus 4 x minus 5 divided by quantity x plus 3?

OpenStudy (anonymous):

Well, this is what i read : R(x) = [(x^2 + 5x + 4) / (x^2 + 3x -4)] . (x^2 + 4x - 5)/(x + 3) now, we have : x^2 + 5x + 4 = (x+1)(x+4) --> check by developing... x^2 + 3x - 4 = (x-1)(x+4) and second factor : x^2 + 4x - 5 = (x-1)(x+5) therefore : R(x) = [ (x+1)(x+4) / [(x-1)(x+4)] ] . [ (x-1)(x+5) / (x+3) ] R(x) = [ (x+1) / (x-1) ] . [ (x-1)(x+5) / (x+3) ] R(x) = (x+1) . (x+5) / (x+3) finally : in terms of restrictions, we can say that : -----> R(x) is defined on : R \ { -3 } hope it' ll help !!